OpenStudy (lncognlto):

To solve a trigonometric equation:

OpenStudy (lncognlto):

\[\sqrt{3}sinx = \frac{ 1 }{ 2 } secx\]

OpenStudy (lncognlto):

\[2\sqrt{3}sinx = secx\]

OpenStudy (lncognlto):

\[2\sqrt{3}sinx - \frac{ 1 }{ cosx } = 0\]

OpenStudy (lncognlto):

Now where do I go?

OpenStudy (anonymous):

now you find for which values of x make \[2\sqrt{3}\sin x=\frac{ 1 }{ cosx}\] true. it's easiest just to look at a unit circle